The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., whatt wioud be the width of the rectangle.
Answer by jai_kos(139)

The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., what would be the width of the rectangle?
Solution:
Given length = 3 + 2 * width (1)
Perimeter = 36 inches. (2)
Formula for perimeter = 2 (length) + 2 (width)
Substitute the value from the equations (1) and (2) in the formula, we get
36 = 2 (3+ 2 * width) + 2*width
36 = 6 + 4 width + 2 width
36 = 6 + 6width
Subtract both sides by 6
36 -6 = 6 + 6width -6
30 = 6 width
Divide the above equation by 6, we get
(30) /6 = width
5 = width.
That is width = 5 inches.